How to derive the price from TR=100,000Q-10Q^2
1
Expert's answer
2015-07-27T13:44:00-0400
TR = P*Q = (100 - 10*Q)*Q .
Then P = 100 - 10*Q .
To maximaze TR we have to find an optimal Q:
MR = (d TR)/(d Q) = 100 - 20Q = 0 .
Then we can find Q = 5 and P = 100 - 10*Q = 100 - 10*5 = 50